OpenStudy (anonymous):

if the volume of a box is 1344 cubic inches how many cubic inches are in one cubic foot

OpenStudy (anonymous):

12 inches = 1 foot 12^3 inches = 1^3 foot = 1728 cubic inches..

OpenStudy (anonymous):

ty ty ty ty

OpenStudy (anonymous):

No problem. if you need more help feel free to ask

OpenStudy (anonymous):

c: